Using values of ΔG°f, calculate ΔG°rxn for the following reaction. Is the reaction product-favored or reactant-favored?

SiCl4(g) + 2 Mg(s) → 2 MgCl2(s) + Si(s)

ΔG°f (kJ/mol)
HgS(s) -50.6
SO2(g) -300.13
H2S(g) -33.56
H2O(ℓ) -228.59
SiCl4(g) -622.76
MgCl2(s) -591.59

____ kJ/mol?

product-favored or reactant-favored?
Answer #1

Reaction taking place is:

SiCl4(g) + 2Mg(s) --> 2MgCl2(s) + Si(s)

dGrxn = dGprod - dGreact

dGprod = 2*dGMgCl2 + dGSi = 2*(-591.59) + 0 = -1183.18 kJ

dGreact = dGSiCl4 + 2*dGMg = -622.76 + 2*0 = -622.76 kJ

So,

dGrxn = -1183.18 - (-622.76) = -560.42 kJ

Since dGrxn < 0, this means the reaction is spontaneous and is hence product favored.

Hope this helps !